The product of Edna's age in two years and her age two years ago is 320. What is Edna's present age?
Answer by stanbon(75887)

The product of Edna's age in two years and her age two years ago is 320. What is Edna's present age?
----------
Equation:
(E+2)(E-2) = 320
-----
E^2 -4 = 320
E^2 = 324
----
Positive solution:
E = sqrt(324) = 18 yrs (Edna's age now)
=================
Cheers,
Stan H.